# Fingerprinting Implementations of Cryptographic Primitives and Protocols that Use Post-Quantum Algorithms

### v3

@@ 2.2 SNARKs
SNARKs involve three components: a generator that creates a Common Reference String (CRS) for shared parameters, a prover that encodes the computation as constraints and generates a proof, and a verifier that validates the proof without evaluating the computation directly. SNARK security is evaluated quantitatively by the computational effort required to forge a false proof, measured in "bits of security". For instance, a SNARK with 40 bits of security means an attacker would need $2^{40}$ operations to forge a proof. Qualitative security is the probability of information leakage, ensuring zero-knowledge under statistical assumptions. For example, a qualitative security level of 20 means that the chance of an adversary successfully breaking the zero-knowledge guarantee (e.g., deducing some information about the private input) is $2^{-20}$. Many SNARKs (e.g., Groth16, PlonK, Marlin, Bulletproofs, Nova) rely on discrete logarithm hardness, which quantum algorithms can efficiently solve, rendering them non-post-quantum secure.
Libraries such as lattice-zksnark[38] are actively working toward implementing PQ SNARKs by utilizing parameters resistant to quantum attacks. For example, lattice-zksnark uses a quantitative security of 128 bits and a qualitative security of 40. Other efforts in this domain include zkLLVM[27], which explores compiling SNARK-friendly circuits with post-quantum backends, and research prototypes based on Starks (Scalable Transparent Arguments of Knowledge), which use hash-based security and do not rely on discrete logarithms. Solana, a high performance blockchain introduced the "Solana Winternitz Vault," a quantum-resistant, hash-based signature system to secure user funds against quantum threats. This optional feature regenerates private keys per transaction, enhancing security for those who opt in.
@@ 2.3 Integration of Post Quantum KEM and Digital Signature in Secure Protocols
The adoption of PQ algorithms in widely used secure protocols, such as TLS, QUIC, VPN, Open Id Connect (OIDC), and SSH, is critical to mitigating future quantum threats. These protocols are transitioning through hybrid approaches, combining classical and PQ cryptography for backward compatibility. For instance, TLS and QUIC are integrating PQ key exchanges, VPNs are employing PQ KEMs for encrypted tunnels, OIDC is updating identity token signing and TLS handshakes with PQ signatures and key exchanges, and SSH is incorporating PQ algorithms in its key exchange processes. This phased approach allows existing systems to maintain security as they gradually shift to post-quantum cryptographic standards. We provide more details in Table 2.
Hybrid Key Exchange in TLS 1.3. Every TLS connection starts with a handshake which is necessary to securely establish a shared encryption key and negotiate encryption protocols before any data is exchanged, ensuring the confidentiality and integrity of the connection. It begins with the ClientHello message, where the client proposes cryptographic settings for key exchange. TLS 1.3 ClientHello messages always contain an extensions field (minimally "supported_versions") which contain the public keys of the algorithms the client supports. This is followed by the ServerHello, where the server agrees on the parameters. The server then sends a Server Certificate and may request a client certificate. Finally, the client and server exchange Finished messages after generating session keys, completing the handshake and allowing secure data transmission. Below we describe how TLS key exchange is changed in the PQ and hybrid scenarios.
Classical. In TLS 1.3, the client initiates the Diffie-Hellman (DH) key exchange by selecting a secret exponent x and computing the corresponding public key $g^{x}$, which is then included in the extension field of the ClientHello. Upon receiving this, the server selects its own secret exponent y and computes its public key $g^{y}$, transmitting it in the ServerHello. The shared secret $g^{xy}$ is derived independently by both the client and server (see Figure 3), enabling secure communication.
@@
Hybrid. In a hybrid key exchange within TLS 1.3, both classical and post-quantum keys are generated. The classical key is derived from methods like Diffie-Hellman (DH) or Elliptic Curve Diffie-Hellman (ECDHE), while the post-quantum key is obtained from a quantum-resistant algorithm. The client’s ClientHello message contains the concatenated post-quantum and classical public keys as an entry of the extensions field. These keys are concatenated without any additional encoding and transmitted as a single value, avoiding alterations to existing protocol data structures. The server then generates a random secret, encrypts it with the client’s post-quantum public key, and sends the ciphertext along with its classical public key in the ServerHello. The client decrypts the ciphertext using its post-quantum private key to retrieve the post-quantum shared secret, while the classical shared secret is computed using the classical key exchange method. The final shared secret is derived by concatenating the results of both the classical and post-quantum exchanges (see Figure 4). This information is sourced from the draft available at [24].
@@
TLS (Transport Layer Security) TLS operates on top of the TCP layer to secure data over networks with encrypted handshakes to authenticate parties and establish secure channels. Post-quantum handshakes replace classical methods with post-quantum or hybrid keys. OpenSSL[71]
Schannel [49]
Secure Transport API[7]
s2n-TLS[12]
OQS-OpenSSL[63]
AWS-LC[6]
@@
QUIC (Quick UDP Internet Connections) QUIC, a transport protocol by Google, enhances communication with 0-RTT connections, TLS 1.3 encryption, multiplexing, and seamless migration. Post-quantum handshakes use post-quantum or hybrid keys instead of classical ones. Quiche[20]
gQUIC[72]
s2n-Quic[11] AWS-LC
VPN (Virtual Private Network) VPNs create secure, encrypted tunnels over the internet, protecting data and ensuring privacy during remote network access or internet use. They use protocols like IPsec or TLS, with TLS securing initial connections. Post-quantum handshakes replace classical keys with post-quantum or hybrid keys to counter quantum threats. OpenVPN[65] Wireguard[23] L2TP/IPSEC[37] IPSEC[36]
PPTP[46]
SSTP[47]
PQ-crypto-OpenVPN[64]
PQ-Wireguard[4]
@@
OIDC (OpenID Connect) OpenID Connect (OIDC), built on OAuth 2.0, enables clients to verify user identities via an authorization server. The OIDC Provider (OP) issues ID tokens, which Relying Parties (RP) use to authenticate users and grant access. Post-quantum TLS handshakes secure communication, while the OP employs post-quantum signatures for signing JWT tokens, protecting against quantum threats. Auth0[10]
Okta[61]
Google Identity[30]
Entra ID[48]
Post-quantum-oidc-oauth2 [68] liboqs
SSH (Secure Shell) SSH is a protocol for securely accessing and managing remote servers via encrypted communication. It enables secure command execution, file transfers, and port forwarding. Post-quantum handshakes replace classical key exchanges with post-quantum or hybrid keys to counter quantum threats. OpenSSH (version < 9.0)[75]
@@
Table 2: Secure Protocols Supporting Post-Quantum Algorithms
3 Fingerprinting of Libraries
In this section, we describe the fingerprinting of key exchange algorithms and signature schemes in library implementations. We first describe our methodology, then we describe our algorithm, then describe the detection for key exchange and digital signatures. We seek to answer the following questions:
@@ 3.1 Methodology
Libraries we consider. We run experiments on publicly available libraries such as liboqs  and CIRCL  that implement PQ key exchange and signature algorithms. We did not utilize pqm4  in our experiments because it is specifically optimized for the ARM Cortex-M4 microcontroller, which is primarily used in embedded systems rather than general-purpose computing environments. Our experiments focused on evaluating PQ cryptographic algorithms within operating systems such as Ubuntu, macOS, and Windows, which are representative of desktop and server environments where the broader deployment of cryptographic protocols is more relevant. We did not consider AWS-LC library because it has only 1 implementation of CRYSTALS-Kyber. Implementations of hybrid algorithms were only available in CIRCL. Classical implementations of algorithms were taken from the libtomcrypt  library. The liboqs  and libtomcrypt  libraries are written in C, while CIRCL  is written in Go.
Available algorithms. Classical key exchange algorithms include RSA[74] and ECDH variants. Classical signature schemes include variants of DSA[69], RSA and ECDSA[41].
@@
Device Specifications Experiments were conducted on three different computing devices. The first device was equipped with Ubuntu 22.04, powered by an Intel Core i5-11400H CPU, which features 6 physical cores and 12 virtual cores, complemented by 16 GB of RAM. The second device operated on Windows 10, utilizing an Intel Core i5-6300U CPU with 2 physical cores and 4 virtual cores, also paired with 16 GB of RAM. The third device, also running Windows 10, was configured with an Intel Core i5-4300U CPU, with 2 physical cores and 4 virtual cores, and 8 GB of RAM.
Datasets. CPU cycle count and memory usage were considered as features for the classification tasks. The libraries have dependencies on system calls unique to Linux systems and so cannot be carried out solely on Windows without WSL[50]. The perf command was used to get a per core cycle count of the running process. For memory usage we captured the memory imprint of a running process from the /proc/[pid]/status file in Linux. The cycle counts and memory usage for the processes were generated across 1000 runs.
@@
3.2 Fingerprinting Method
We formulate the fingerprinting problem as a machine learning classification, with classes identified by the names of the cryptographic algorithms. For each classification scenario in key exchange, we start the classification task with 19 features of which 12 features represent the cycles used by each core of the CPU and remaining 7 features represent a process’s memory usage, including total virtual memory(VMSize), physical memory in use(VMRSS), data segment(VMData), stack, executable code(VMStk), shared libraries(VMLib), and page table entries(VMPTE). For devices having less than 12 virtual CPU cores the additional fields were kept empty and we did not have devices with more than 12 cores. We narrow down our feature set to 4 memory based features (VMSize, VMRSS, VMData, VMExe) using the Chi-square statistical test as the scoring function to get the top features by importance. In digital signatures, we used the same 4 features in our set for two classification scenarios but for the multi-class scenario we needed all 19 features. We do an 80-20 train-test split of the dataset and fit models such as Logistic Regression, Random Forest, MLP and XGBoost to the data.
@@ 3.3 Key Exchange Results
This section presents our findings across four classification scenarios, which form the basis for answering to questions related to the fingerprinting of PQ key exchange algorithm implementations in libraries. We begin with the classification between classical and PQ algorithms, followed by a multi-class classification focused exclusively on PQ algorithms. Next, we distinguish between the two libraries used in our experiments based on their algorithm implementations. Finally, we demonstrate the binary classification between PQ and hybrid algorithms.
RQ1: Classical vs PQ classification. The datasets consisted of 45,900 samples, with 24,000 classical and 21,900 PQ. Ensemble models, particularly Random Forest and XGBoost, showed superior performance, achieving 98% overall accuracy in Dataset2, with a further increase to 100% in Dataset1. These results indicate that classical and PQ algorithms are readily distinguishable based on memory footprint alone. Detailed metrics are provided in Tables 3 and 4.
@@
Table 3: Key exchange, classical vs PQ, Dataset1
@@
Table 4: Key exchange, classical vs PQ, Dataset2
RQ2: Multi-class PQ algorithm classification. The datasets comprised 21,000 samples, including 3,000 samples from each PQ algorithm (BIKE, SIKE, CRYSTALS-Kyber, FRODO, Classic-McEliece, HQC, NTRUPrime), with equal representation of all variants. Ensemble learning models, particularly Random Forest and XGBoost, outperformed others, achieving 97% and 96% accuracy respectively and strong metrics in identifying individual algorithms within Dataset2. In Dataset1, these models exhibited even higher performance, with 100% accuracy. These results indicate significant differences in memory usage across post-quantum implementations. Detailed results are presented in Tables 5 and 6.
@@
Table 5: Key exchange, PQ algorithms, Dataset1
@@
Table 6: Key exchange, PQ algorithms, Dataset2
RQ3: liboqs  vs CIRCL  classification for FRODO and Kyber. FRODO and CRYSTALS-Kyber were the only algorithms shared between the two libraries, resulting in a dataset of 12,000 samples — 3,000 from each algorithm’s liboqs  and CIRCL  implementations, with equal representation across all variants. CIRCL’s implementation includes one variant of FRODO and all variants of CRYSTALS-Kyber, while liboqs  supports multiple FRODO variants. Random Forest and XGBoost achieved overall accuracy of 96% in Dataset2, rising to 100% in Dataset1. These results demonstrate that libraries are readily distinguishable due to significant differences in cycle counts and memory footprints, influenced by the distinct languages used in their implementations. Detailed metrics are shown in Tables 7 and 8.
@@
Table 7: Key exchange, liboqs vs CIRCL, Dataset1
RQ4: PQ vs Hybrid classification for CIRCL. CIRCL  was the only library supporting both hybrid and post-quantum algorithms. The datasets comprised 29,998 samples: 15,000 for basic post-quantum algorithms (FRODO, CRYSTALS-Kyber, SIKE) and 14,998 for hybrid algorithms. All models achieved 98% accuracy for Dataset1 and 97% overall accuracy for Dataset2. These results indicate that post-quantum and hybrid implementations are distinguishable, likely due to the increased memory usage and computational cost associated with hybrid approaches. Detailed metrics are provided in Tables 9 and 10.
@@
Table 8: Key exchange, liboqs vs CIRCL, Dataset2
@@
Table 9: Key exchange, PQ vs Hybrid, Dataset1
@@
Table 10: Key exchange, PQ vs Hybrid, Dataset2

### v4

@@ 2.2 SNARKs
SNARKs involve three components: a generator that creates a Common Reference String (CRS) for shared parameters, a prover that encodes the computation as constraints and generates a proof, and a verifier that validates the proof without evaluating the computation directly. SNARK security is evaluated quantitatively by the computational effort required to forge a false proof, measured in “bits of security”. For instance, a SNARK with 40 bits of security means an attacker would need $2^{40}$ operations to forge a proof. Qualitative security is the probability of information leakage, ensuring zero-knowledge under statistical assumptions. For example, a qualitative security level of 20 means that the chance of an adversary successfully breaking the zero-knowledge guarantee (e.g., deducing some information about the private input) is $2^{-20}$. Many SNARKs (e.g., Groth16, PlonK, Marlin, Bulletproofs, Nova) rely on discrete logarithm hardness, which quantum algorithms can efficiently solve, rendering them non-post-quantum secure.
Libraries such as lattice-zksnark[ISW21] are actively working toward implementing PQ SNARKs by utilizing parameters resistant to quantum attacks. For example, lattice-zksnark uses a quantitative security of 128 bits and a qualitative security of 40. Other efforts in this domain include zkLLVM[zkllvm], which explores compiling SNARK-friendly circuits with post-quantum backends, and research prototypes based on STARKs (Scalable Transparent Arguments of Knowledge), which use hash-based security and do not rely on discrete logarithms. Solana, a high performance blockchain introduced the “Solana Winternitz Vault”, a quantum-resistant, hash-based signature system to secure user funds against quantum threats. This optional feature regenerates private keys per transaction, enhancing security for those who opt in.
@@ 2.3 Integration of Post Quantum KEM and Digital Signature in Secure Protocols
The adoption of PQ algorithms in widely used secure protocols, such as TLS, QUIC, VPN, Open Id Connect (OIDC), and SSH, is critical to mitigating future quantum threats. These protocols are transitioning through hybrid approaches, combining classical and PQ cryptography for backward compatibility. For instance, TLS and QUIC are integrating PQ key exchanges, VPNs are employing PQ KEMs for encrypted tunnels, OIDC is updating identity token signing and TLS handshakes with PQ signatures and key exchanges, and SSH is incorporating PQ algorithms in its key exchange processes. This phased approach allows existing systems to maintain security as they gradually shift to post-quantum cryptographic standards. We provide more details in Table II.
Hybrid Key Exchange in TLS 1.3. Every TLS connection starts with a handshake which is necessary to securely establish a shared encryption key and negotiate encryption protocols before any data is exchanged, ensuring the confidentiality and integrity of the connection. It begins with the ClientHello message, where the client proposes cryptographic settings for key exchange. TLS 1.3 ClientHello messages always include an extensions field (minimally “supported_versions”) which contains the public keys of the algorithms the client supports. This is followed by the ServerHello message, where the server agrees on the parameters. The server then sends a Server Certificate and may request a client certificate. Finally, the client and server exchange Finished messages after generating session keys, completing the handshake and allowing secure data transmission. Below we describe how TLS key exchange is changed in the PQ and hybrid scenarios.
Classical. In TLS 1.3, the client initiates the Diffie-Hellman key exchange by selecting a secret key x and computing the corresponding public key $g^{x}$, which is then included in the extension field of the ClientHello. Upon receiving this, the server selects its own secret key y and computes its public key $g^{y}$, transmitting it in the ServerHello. The shared secret $g^{xy}$ is derived independently by both the client and server, enabling secure communication (see Figure 1).
@@
Hybrid. In a hybrid key exchange within TLS 1.3, both classical and post-quantum keys are generated. The classical key is derived from methods like Diffie-Hellman (DH) or Elliptic Curve Diffie-Hellman (ECDHE), while the post-quantum key is obtained from a quantum-resistant algorithm. The client’s ClientHello message contains the concatenated post-quantum and classical public keys as an entry of the extensions field. These keys are concatenated without any additional encoding and transmitted as a single value, avoiding alterations to existing protocol data structures. The server then generates a random secret, encrypts it with the client’s post-quantum public key, and sends the ciphertext along with its classical public key in the ServerHello. The client decrypts the ciphertext using its post-quantum private key to retrieve the post-quantum shared secret, while the classical shared secret is computed using the classical key exchange method. The final shared secret is derived by concatenating the results of both the classical and post-quantum exchanges (see Figure 2). This information is sourced from the draft available at [draft-ietf-tls-hybrid-design-10].
Figure 1: Simplified Classical Key Exchange in TLS 1.3
Figure 2: Hybrid Key Exchange in TLS 1.3
@@
TLS (Transport Layer Security) TLS operates on top of the TCP layer to secure data over networks with encrypted handshakes to authenticate parties and establish secure channels. Post-quantum handshakes replace classical methods with post-quantum or hybrid keys. OpenSSL[openssl]
Schannel [schannel]
Secure Transport API[secure-transport]
s2n-TLS[s2n-tls]
OQS-OpenSSL[oqs-openssl]
wolfSSL[wolfssl_pqc]
AWS-LC[aws-lc]
@@
wolfCrypt
QUIC (Quick UDP Internet Connections) QUIC, a transport protocol by Google, enhances communication with 0-RTT connections, TLS 1.3 encryption, multiplexing, and seamless migration. Post-quantum handshakes use post-quantum or hybrid keys instead of classical ones. Quiche[quiche]
gQUIC[gquic]
s2n-Quic[s2n-quic] AWS-LC
VPN (Virtual Private Network) VPNs create secure, encrypted tunnels over the internet, protecting data and ensuring privacy during remote network access or internet use. They use protocols like IPsec or TLS, with TLS securing initial connections. Post-quantum handshakes replace classical keys with post-quantum or hybrid keys to counter quantum threats. OpenVPN[openvpn] Wireguard[wireguard] L2TP/IPSEC[l2tp-ipsec] IPSEC[ipsec]
PPTP[pptp]
SSTP[sstp]
PQ-crypto-OpenVPN[pq-openvpn]
PQ-Wireguard[pq-wireguard]
@@
OIDC (OpenID Connect) OpenID Connect (OIDC), built on OAuth 2.0, enables clients to verify user identities via an authorization server. The OIDC Provider (OP) issues ID tokens, which Relying Parties (RP) use to authenticate users and grant access. Post-quantum TLS handshakes secure communication, while the OP employs post-quantum signatures for signing JWT tokens, protecting against quantum threats. Auth0[auth0]
Okta[okta]
Google Identity[google-identity]
Entra ID[entra-id]
Post-quantum-oidc-oauth2 [postquantum-oidc-oauth2] liboqs
SSH (Secure Shell) SSH is a protocol for securely accessing and managing remote servers via encrypted communication. It enables secure command execution, file transfers, and port forwarding. Post-quantum handshakes replace classical key exchanges with post-quantum or hybrid keys to counter quantum threats. OpenSSH (version ¡ 9.0)[openssh]
@@
TABLE II: Secure Protocols Supporting Post-Quantum Algorithms
3 Classification of Libraries
In this section, we describe the classification of key exchange algorithms and signature schemes in library implementations. We first describe our methodology, then we describe our algorithm, then describe the detection for key exchange and digital signatures. We seek to answer the following questions:
@@ 3.1 Methodology
Libraries we consider. We run experiments on publicly available libraries such as liboqs  and CIRCL  that implement PQ key exchange and signature algorithms. We did not utilize pqm4  in our experiments because it is specifically optimized for the ARM Cortex-M4 microcontroller, which is primarily used in embedded systems rather than general-purpose computing environments. Our experiments focused on evaluating PQ cryptographic algorithms within operating systems such as Ubuntu, macOS, and Windows, which are representative of desktop and server environments where the broader deployment of cryptographic protocols is more relevant. Integrating pqm4  alongside the other libraries would yield results that are not directly comparable or representative of the broader deployment contexts of cryptographic protocols. We did not consider AWS-LC library because it has only 1 implementation of CRYSTALS-Kyber. The PQClean  project, although providing verified and portable reference implementations, covers only a limited subset of PQ algorithms relative to liboqs  and CIRCL. Incorporating it would have introduced sampling bias toward that restricted subset. We also attempted to compile libpqcrypto, but the build process consistently failed due to compile-time errors, preventing its inclusion in the benchmark suite. Implementations of hybrid algorithms were only available in CIRCL. Classical implementations of algorithms were taken from the libtomcrypt  library. The liboqs  and libtomcrypt  libraries are written in C, while CIRCL  is written in Go.
Available algorithms. Classical key exchange algorithms include RSA[rsa1978] and ECDH variants. Classical signature schemes include variants of DSA[rfc6979], RSA and ECDSA[ecdsa].
@@
Device specifications. Experiments were conducted on three different computing devices. The first device was equipped with Ubuntu 22.04, powered by an Intel Core i5-11400H CPU, which features 6 physical cores and 12 virtual cores, complemented by 16 GB of RAM. The second device operated on Windows 10, utilizing an Intel Core i5-6300U CPU with 2 physical cores and 4 virtual cores, also paired with 16 GB of RAM. The third device, also running Windows 10, was configured with an Intel Core i5-4300U CPU, with 2 physical cores and 4 virtual cores, and 8 GB of RAM.
Datasets. CPU cycle count and memory usage were considered as features for the classification tasks. The libraries have dependencies on system calls unique to Linux systems and so cannot be carried out solely on Windows without WSL[wsl]. The perf command was used to get a per core cycle count of the running process. Although perf inherently reports system-wide cycle counts, background activity was minimized to ensure that the reported measurements provided a sufficiently accurate approximation of the per-core cycle count attributable to the running process. For memory usage we captured the memory imprint of a running process from the /proc/[pid]/status file in Linux. The cycle counts and memory usage for the processes were generated across 1000 runs.
@@
3.2 Classification Method
We formulate the problem as a machine learning classification, with classes identified by the names of the cryptographic algorithms. For each classification scenario in key exchange, we start the classification task with 19 features of which 12 features represent the cycles used by each core of the CPU and remaining 7 features represent a process’s memory usage, including total virtual memory(VMSize), physical memory in use(VMRSS), data segment(VMData), stack, executable code(VMStk), shared libraries(VMLib), and page table entries(VMPTE). For devices having less than 12 virtual CPU cores the additional fields were kept empty and we did not have devices with more than 12 cores. We narrow down our feature set to 4 memory based features (VMSize, VMRSS, VMData, VMExe) using the Chi-square statistical test as the scoring function to get the top features by importance. In digital signatures, we used the same 4 features in our set for two classification scenarios but for the multi-class scenario we needed all 19 features. We do an 80-20 train-test split of the dataset and fit models such as Logistic Regression, Random Forest, MLP and XGBoost to the data.
@@ 3.3 Key Exchange Results
This section presents our findings across four classification scenarios, which form the basis for answering to questions related to the classification of PQ key exchange algorithm implementations in libraries. We begin with the classification between classical and PQ algorithms, followed by a multi-class classification focused exclusively on PQ algorithms. Next, we distinguish between the two libraries used in our experiments based on their algorithm implementations. Finally, we demonstrate the binary classification between PQ and hybrid algorithms.
RQ1: Classical vs PQ classification. The datasets consisted of 45,900 samples, with 24,000 classical and 21,900 PQ. Ensemble models, particularly Random Forest and XGBoost, showed superior performance, achieving 98% overall accuracy in Dataset2, with a further increase to 100% in Dataset1. These results indicate that classical and PQ algorithms are readily distinguishable based on memory footprint alone. Detailed metrics are provided in Tables III and IV.
@@
TABLE III: Key exchange, classical vs PQ, Dataset1
@@
TABLE IV: Key exchange, classical vs PQ, Dataset2
RQ2: Multi-class PQ algorithm classification. The datasets comprised 21,000 samples, including 3,000 samples from each PQ algorithm (BIKE, SIKE, CRYSTALS-Kyber, FRODO, Classic-McEliece, HQC, NTRUPrime), with equal representation of all variants. Ensemble learning models, particularly Random Forest and XGBoost, outperformed others, achieving 97% and 96% accuracy respectively and strong metrics in identifying individual algorithms within Dataset2. In Dataset1, these models exhibited even higher performance, with 100% accuracy. These results indicate significant differences in memory usage across post-quantum implementations. Detailed results are presented in Tables V and VI.
@@
TABLE V: Key exchange, PQ algorithms, Dataset1
@@
TABLE VI: Key exchange, PQ algorithms, Dataset2
RQ3: liboqs  vs CIRCL  classification for FRODO and Kyber. FRODO and CRYSTALS-Kyber were the only algorithms shared between the two libraries, resulting in a dataset of 12,000 samples — 3,000 from each algorithm’s liboqs  and CIRCL  implementations, with equal representation across all variants. CIRCL’s implementation includes one variant of FRODO and all variants of CRYSTALS-Kyber, while liboqs  supports multiple FRODO variants. Random Forest and XGBoost achieved overall accuracy of 96% in Dataset2, rising to 100% in Dataset1. These results demonstrate that libraries are readily distinguishable due to significant differences in cycle counts and memory footprints, influenced by the distinct languages used in their implementations. Detailed metrics are shown in Tables VII and VIII.
@@
TABLE VII: Key exchange, liboqs vs CIRCL, Dataset1
RQ4: PQ vs Hybrid classification for CIRCL. CIRCL  was the only library supporting both hybrid and post-quantum algorithms. The datasets comprised 29,998 samples: 15,000 for basic post-quantum algorithms (FRODO, CRYSTALS-Kyber, SIKE) and 14,998 for hybrid algorithms. All models achieved 98% accuracy for Dataset1 and 97% overall accuracy for Dataset2. These results indicate that post-quantum and hybrid implementations are distinguishable, likely due to the increased memory usage and computational cost associated with hybrid approaches. Detailed metrics are provided in Tables IX and X.
@@
TABLE VIII: Key exchange, liboqs vs CIRCL, Dataset2
@@
TABLE IX: Key exchange, PQ vs Hybrid, Dataset1
@@
TABLE X: Key exchange, PQ vs Hybrid, Dataset2